OpenStudy (anonymous):

y = x^2 + 10x + 20 write the equation in vertex form. identify the vertex, axis of symmetry and directon of opening

OpenStudy (freckles):

first of all do you know how to complete the square?

OpenStudy (anonymous):

no

OpenStudy (freckles):

\[\[x^2+kx+(\frac{k}{2})^2 \\ \text{ this can be written as } (x+\frac{k}{2})^2\] \]

OpenStudy (freckles):

\[y=(x^2+10x)+20 \\ y=(x^2+10x+?)+20-?\]

OpenStudy (freckles):

so what can you replace that question with so that the thing in ( ) can be written as something to the second power?

OpenStudy (anonymous):

I had originally simplified it to y=x(x+10)+20. I'm not sure how to figure out what y would equal

OpenStudy (anonymous):

I know that doesn't answer your question, but I'm not sure

OpenStudy (freckles):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2 \\ x^2+10x+(\frac{10}{2})^2=(x+\frac{10}{2})^2\]

OpenStudy (freckles):

so looking back at \[y=(x^2+10x+?)+20-?\] the ? needs to be what

OpenStudy (freckles):

I just gave you a big hint

OpenStudy (freckles):

we are trying to add something in to complete the square I actually already told you what to add in you just have to tell me what it was I told you to add in

OpenStudy (freckles):

and of course whatever we add in we must subtract out

OpenStudy (freckles):

that is why there is +?-?

OpenStudy (anonymous):

right but that's where I'm stuck. I don't get that.

OpenStudy (freckles):

here is a general example when the coefficient of x^2 is 1: \[y=x^2+bx+c\] \[y=(x^2+bx)+c \\ y=(x^2+bx+(\frac{b}{2})^2)+c-(\frac{b}{2})^2 \\ y=(x+\frac{b}{2})^2+c-(\frac{b}{2})^2\]